# MM-Spatial: Exploring 3D Spatial Understanding in Multimodal LLMs

### v1

@@ Abstract
Multimodal large language models (MLLMs) excel at 2D visual understanding but remain limited in their ability to reason about 3D space. In this work, we leverage large-scale high-quality 3D scene data with open-set annotations to introduce 1) a novel supervised fine-tuning dataset and 2) a new evaluation benchmark, focused on indoor scenes. Our Cubify Anything VQA (CA-VQA) data covers diverse spatial tasks including spatial relationship prediction, metric size and distance estimation, and 3D grounding. We show that CA-VQA enables us to train MM-Spatial, a strong generalist MLLM that also achieves state-of-the-art performance on 3D spatial understanding benchmarks, including our own. We show how incorporating metric depth and multi-view inputs (provided in CA-VQA) can further improve 3D understanding, and demonstrate that data alone allows our model to achieve depth perception capabilities comparable to dedicated monocular depth estimation models. We will publish our SFT dataset and benchmark. MM-Spatial: Exploring 3D Spatial Understanding in Multimodal LLMs
@@ 1 Introduction
Understanding object locations and spatial relationships in both 2D and 3D space is crucial for interpreting complex visual scenes. While multimodal large language models (MLLMs) have achieved notable success for 2D visual tasks including referring and grounding [122, 127] and spatial relation prediction (e.g., ”left” vs. ”right”, ”above” vs. ”below”) [27], they still struggle with 3D object perception tasks such as estimating 1) relative depth (”in front” vs. ”behind”), 2) object distances or sizes in metric units (”A is 2.74m away / 1.32m wide.”), and, ultimately, 3) precise 3D bounding boxes. Yet, the ability to reason about objects in 3D scenes is not only a part of general visual comprehension, but is also foundational in domains like robotics and AR / VR, e.g., for navigation and manipulation tasks [83].
Dataset Data source(s) High-quality 3D Ground-truth Depth maps Multi-view images Tasks Splits Public
@@ 2.1 General MLLMs
MLLMs research has explored several fronts: ($i$) scaling up the pre-training data [76, 85, 115, 9, 68] and supervised fine-tuning data [48, 102, 60, 104]; ($ii$) enhancing high-resolution image comprehension [78, 71, 81, 31, 40, 41, 24, 131, 114, 70]; ($iii$) studying various vision encoders [105, 96, 22, 36] and vision-language connectors [17, 118, 69, 14]; ($iv$) using mixture-of-experts [74, 66]; ($v$) extending models to region-level [110, 133, 125, 88, 23, 130, 122, 127] and pixel-level [58, 91, 124, 94] understanding, multi-image reasoning [52, 65], UI comprehension [123, 72, 45], and video understanding [75, 112, 113, 120], among others.
@@ 3.1 Data Generation Pipeline
  • QA Pair Generation. Based on our 3D and semantic annotations, we automatically generate template-based QA pairs (both open-ended and multi-choice ones), without any human supervision. We consider a variety of spatial task categories, detailed in Sec. 3.2 and App. A. We further ensure that questions are unambiguous. For example, asking ”What is the distance to the chair?”, is only a valid question is there is a single instance of a chair.
@@ 3.2 Spatial Task Categories
Counting (”How many X are there?”). Answers are computed by counting the 3D boxes of the given object class.
Viewpoint-dependent (”Is X behind Y?”). Answers are based on the 2D / 3D boxes and depend on the camera pose.
Metric regression (”How far away is X from Y / the camera?” ”How wide / tall is X?”) Answers to size questions are computed using the 3D bounding box. Answers to distance questions are computed based on the object point clouds; we reject samples for which the 3D boxes overlap.
@@
Binary (e.g., ”Is X taller than Y?”). This covers viewpoint-dependent and (relative) regression questions, as well as object presence questions (”Is X present in the image?”).
@@ 4 Model
[…]
Figure 4: Qualitative Example. We show the predictions of various models on a challenging example from our CA-VQA benchmark. Strong commercial (2a&b) and research models (2c&d) fail. MM-Spatial (1a) is much better, and even more so with CoT enabled (1b), demonstrating our model’s strong object grounding (see predicted 2D boxes in the image), depth estimation, and spatial reasoning ability. Accuracy improves further when leveraging ground-truth depth via tool-use (1c), although our CoT model’s (1b) predictions are very close to that, for both the intermediate depth values and final answer; monocular estimated depth (1d) is less accurate and yields a worse result.
@@ 5.3 Results on our CA-VQA Benchmark
  • MM-Spatial vs. Baselines. MM-Spatial-3B 8 substantially outperforms various (much larger) top open-source and commercial models 1- 6, incl. the SOTA GPT-4o model 3, demonstrating 1) their limitations in terms of spatial understanding and 2) the effectiveness of SFT on CA-VQA. Despite its focus on spatial reasoning, SpatialRGPT-VILA-1.5-8B 6 underperforms, likely for a few reasons: 1) their OpenSpatialDataset (OSD) used for SFT leverages axis-aligned 3D boxes pseudo-annotated on OpenImages [56], resulting in significant discrepancies in spatial concept definitions (especially for metric quantities) compared to the high-quality (yaw-)oriented 3D boxes used for CA-VQA; see App. E for further discussion on this difference; 2) in OSD, objects are referred to via segmentation masks or 2D bounding boxes (”How tall is Region [0] <mask/box>?”), while CA-VQA simply uses class names (”How tall is the chair?”); 3) SpatialRGPT is limited to using relative depth due to its full depth encoding approach. Finally, SpatialRGPT / OSD lacks support for 3D grounding.
@@
    Model ${}^{\text{2D}}$) ${}^{\text{3D}}$) Average (2D+3D)
    Object Count Spatial Relation. Average (2D) Depth Order Relative Distance Average (3D)
@@
Below / Above Left / Right Big / Small Tall / Short Wide / Thin Behind / Front Avg. Direct Dist. Horizon. Dist. Vertical Dist. Width Height Avg.
@@
7 MM-Spatial-3B CA-VQA (our defs.) 98.3 97.1 80.2 82.1 73.1 74.6 84.2 14.2 12.3 47.2 30.0 53.4 31.4 57.8
8 MM-Spatial-3B CA-VQA + OSD 98.3 99.1 94.3 93.8 92.3 93.6 95.2 33.1 74.6 61.3 55.6 77.4 60.4 77.8
@@ 5.4 CV-Bench Results
The CV-Bench results in Tab. 4 demonstrate that MM-Spatial-3B 10 significantly outperforms the much larger SOTA Cambrian-1-34B 8, highlighting the effectiveness of SFT on similar data. CoT 11 and leveraging monocular (DepthPro) depth input via tool-use 12 again further boost performance. MM-Spatial achieves almost perfect accuracy on the indoor splits of the 3D tasks, and also demonstrates strong out-of-domain generalization to the outdoor splits. Notably, MM-Spatial (Blind eval) 13 achieves the best accuracy among all models on the 2D Object Count task, revealing a substantial bias in this benchmark. In contrast, on our CA-VQA benchmark, using vision input outperforms the blind baseline on Counting by $\sim$13 points.
@@ 5.5 SpatialRGPT-Bench Results
  • CA-VQA vs. OSD. Training on OSD 20 (based on diverse OpenImages [56] data) results in strong performance both indoor 21 and outdoor 22. However, training on CA-VQA still yields significantly better indoor performance overall 15, suggesting that our high-quality 3D GT is more effective than OSD’s pseudo-annotations. When using CA-VQA + scale aug. 19, we become competitive with OSD 22 even on outdoor Distance, but still lack on outdoor Width / Height. Combining CA-VQA with OSD 13 results in significant improvements, emphasizing the complementary benefits of the two datasets.
@@ A.1 Spatial Task Categories
    • Viewpoint-Dependent. We consider the spatial relationships left vs. right and in front vs. behind between two objects, as determined from the current camera pose / viewpoint: [^2] Note that we do not consider above vs. below to avoid ambiguity: ”above” could either refer to 2D image space (i.e., the 2D bounding box of A is above that of B), or to 3D space, where the latter can be ambiguous as well (i.e., do we just require that the 3D bounding box of A is located higher in terms of vertical dimension, or do we also require that A is located directly above B in terms of horizontal dimensions – the latter might best match with how humans colloquially define ”above”).
@@
    • Counting. We always ensure that 0 is an option (i.e., object is not present). We then randomly sample (additional) wrong options among the non-zero integers within $[\text{GT}-3,\text{GT}+3]$ (where GT is the correct answer), s.t. the total number of options is 4.
@@ A.2 Depth: Chain-of-Thought (CoT) / Tool-Use
We prepare multi-step CoT responses involving depth for questions within the Binary (only ”behind vs. in front”) and Regression (Metric Estimation) categories, as the ground truth answers for those rely on depth information. We also did preliminary experiments with 3D Grounding samples, but found that performance does not improve / even slightly regresses there, so we did not include any such samples in the final dataset. [^3] We hypothesize that 3D grounding is too complex of a task to benefit from the simple depth information provided in the multi-step CoT answers, and that the model might just get confused. We leave a more comprehensive study of how to benefit 3D grounding with CoT for future work.
@@ Appendix D Analysis of Blind Filtering Procedure
Table 11: CA-VQA Blind Filtering Analysis. We study how the improvement from using vision (i.e., comparing a vision-evaluated model vs. a blind-evaluated model) changes after applying the blind filtering strategy outlined in Sec. 3.1, which follows [25]. Our results confirm that after applying our filtering strategy, 1) blind models perform substantially worse, and 2) vision improvements (i.e., the delta between vision and blind models) increase substantially, for both GPT-4/V and MM-Spatial. This highlights the effectiveness of our blind filtering procedure in ensuring that our CA-VQA benchmark becomes more reliant on vision input (i.e., less susceptible to a strong language prior).
@@
Figure 8: Comparative visualization of axis-aligned vs. oriented 3D bounding boxes, taken from the SpatialRGPT paper [27, Appendix K, Figure 11]. The object dimensions computed from AABBs can differ substantially from those computed from OBBs, depending on the object’s rotation. For sake of illustration, assume that the sofa is 2m wide and 0.8m deep. We then obtain the following altered object dimensions when using an AABB instead of an OBB, at different yaw rotation angles (i.e., considering 7-DOF bounding boxes that are gravity-aligned / parallel to the ground, as in CA-1M / CA-VQA): width $\approx$ 2.1m and depth $\approx$ 1.7m with 30 rotation; width $\approx$ 1.7m and depth $\approx$ 2.1m with 60 rotation; and width = 0.8m and depth = 2m with 90 rotation (i.e., ”full” rotation resulting in swapped dimensions).

### v2

@@ Abstract
Multimodal large language models (MLLMs) excel at 2D visual understanding but remain limited in their ability to reason about 3D space. In this work, we leverage large-scale high-quality 3D scene data with open-set annotations to introduce 1) a novel supervised fine-tuning dataset and 2) a new evaluation benchmark, focused on indoor scenes. Our Cubify Anything VQA (CA-VQA) data covers diverse spatial tasks including spatial relationship prediction, metric size and distance estimation, and 3D grounding. We show that CA-VQA enables us to train MM-Spatial, a strong generalist MLLM that also achieves state-of-the-art performance on 3D spatial understanding benchmarks, including our own. We show how incorporating metric depth and multi-view inputs (provided in CA-VQA) can further improve 3D understanding, and demonstrate that data alone allows our model to achieve depth perception capabilities comparable to dedicated monocular depth estimation models. https://github.com/apple/ml-cubifyanything
MM-Spatial: Exploring 3D Spatial Understanding in Multimodal LLMs
@@ 1 Introduction
Understanding object locations and spatial relationships in both 2D and 3D space is crucial for interpreting complex visual scenes. While multimodal large language models (MLLMs) have achieved notable success for 2D visual tasks including referring and grounding [122, 127] and spatial relation prediction (e.g., “left” vs. “right”, “above” vs. “below”) [27], they still struggle with 3D object perception tasks such as estimating 1) relative depth (“in front” vs. “behind”), 2) object distances or sizes in metric units (“A is 2.74m away / 1.32m wide.”), and, ultimately, 3) precise 3D bounding boxes. Yet, the ability to reason about objects in 3D scenes is not only a part of general visual comprehension, but is also foundational in domains like robotics and AR / VR, e.g., for navigation and manipulation tasks [83].
Dataset Data source(s)
High-quality
3D Ground-truth
Depth maps
Multi-view
images
Tasks Splits Public
@@ 2.1 General MLLMs
MLLMs research has explored several fronts: ($i$) scaling up the pre-training data [76, 85, 115, 9, 68] and supervised fine-tuning data [48, 102, 60, 104]; ($ii$) enhancing high-resolution image comprehension [78, 71, 81, 31, 40, 41, 24, 131, 114, 70]; (iiiiii) studying various vision encoders [105, 96, 22, 36] and vision-language connectors [17, 118, 69, 14]; ($iv$) using mixture-of-experts [74, 66]; ($v$) extending models to region-level [110, 133, 125, 88, 23, 130, 122, 127] and pixel-level [58, 91, 124, 94] understanding, multi-image reasoning [52, 65], UI comprehension [123, 72, 45], and video understanding [75, 112, 113, 120], among others.
@@ 3.1 Data Generation Pipeline
  • QA Pair Generation. Based on our 3D and semantic annotations, we automatically generate template-based QA pairs (both open-ended and multi-choice ones), without any human supervision. We consider a variety of spatial task categories, detailed in Sec. 3.2 and App. A. We further ensure that questions are unambiguous. For example, asking “What is the distance to the chair?”, is only a valid question is there is a single instance of a chair.
@@ 3.2 Spatial Task Categories
Counting (“How many X are there?”). Answers are computed by counting the 3D boxes of the given object class.
Viewpoint-dependent (“Is X behind Y?”). Answers are based on the 2D / 3D boxes and depend on the camera pose.
Metric regression (“How far away is X from Y / the camera?” “How wide / tall is X?”) Answers to size questions are computed using the 3D bounding box. Answers to distance questions are computed based on the object point clouds; we reject samples for which the 3D boxes overlap.
@@
Binary (e.g., “Is X taller than Y?”). This covers viewpoint-dependent and (relative) regression questions, as well as object presence questions (“Is X present in the image?”).
@@ 4 Model
Figure 3: Example of leveraging depth maps via tool-use. The model predicts the objects’ 2D bounding boxes and function calls, receives the tool outputs (which is the median depth value within the box, marked with an $\mathbf{\times}$), and finally reasons about the answer.
Figure 4: Qualitative Example. We show the predictions of various models on a challenging example from our CA-VQA benchmark. Strong commercial (2a&b) and research models (2c&d) fail. $M$-Spatial (1a) is much better, and even more so with CoT enabled (1b), demonstrating our model’s strong object grounding (see predicted 2D boxes in the image), depth estimation, and spatial reasoning ability. Accuracy improves further when leveraging ground-truth depth via tool-use (1c), although our CoT model’s (1b) predictions are very close to that, for both the intermediate depth values and final answer; monocular estimated depth (1d) is less accurate and yields a worse result.
@@ 5.3 Results on our CA-VQA Benchmark
  • MM-Spatial vs. Baselines. MM-Spatial-3B 8 substantially outperforms various (much larger) top open-source and commercial models 1- 6, incl. the SOTA GPT-4o model 3, demonstrating 1) their limitations in terms of spatial understanding and 2) the effectiveness of SFT on CA-VQA. Despite its focus on spatial reasoning, SpatialRGPT-VILA-1.5-8B 6 underperforms, likely for a few reasons: 1) their OpenSpatialDataset (OSD) used for SFT leverages axis-aligned 3D boxes pseudo-annotated on OpenImages [56], resulting in significant discrepancies in spatial concept definitions (especially for metric quantities) compared to the high-quality (yaw-)oriented 3D boxes used for CA-VQA; see App. E for further discussion on this difference; 2) in OSD, objects are referred to via segmentation masks or 2D bounding boxes (“How tall is Region [0] <mask/box>?”), while CA-VQA simply uses class names (“How tall is the chair?”); 3) SpatialRGPT is limited to using relative depth due to its full depth encoding approach. Finally, SpatialRGPT / OSD lacks support for 3D grounding.
@@
[…]
    Average
    (2D+3D)
    Object
    Count
    Spatial
    Relation.
    Average
    Depth Order Relative Distance
    Average
@@
Below /
Above
Left /
Right
Big /
Small
Tall /
Short
Wide /
Thin
Behind
/ Front
Avg.
Direct
Dist.
Horizon.
Dist.
Vertical
Dist.
Width Height Avg.
@@
7 $M$-Spatial-3B CA-VQA (our defs.) 98.3 97.1 80.2 82.1 73.1 74.6 84.2 14.2 12.3 47.2 30.0 53.4 31.4 57.8
8 $M$-Spatial-3B CA-VQA + OSD 98.3 99.1 94.3 93.8 92.3 93.6 95.2 33.1 74.6 61.3 55.6 77.4 60.4 77.8
@@ 5.4 CV-Bench Results
The CV-Bench results in Tab. 4 demonstrate that MM-Spatial-3B 10 significantly outperforms the much larger SOTA Cambrian-1-34B 8, highlighting the effectiveness of SFT on similar data. CoT $1$ and leveraging monocular (DepthPro) depth input via tool-use 12 again further boost performance. MM-Spatial achieves almost perfect accuracy on the indoor splits of the 3D tasks, and also demonstrates strong out-of-domain generalization to the outdoor splits. Notably, MM-Spatial (Blind eval) 13 achieves the best accuracy among all models on the 2D Object Count task, revealing a substantial bias in this benchmark. In contrast, on our CA-VQA benchmark, using vision input outperforms the blind baseline on Counting by $\sim$13 points.
@@ 5.5 SpatialRGPT-Bench Results
  • CA-VQA vs. OSD. Training on OSD 20 (based on diverse OpenImages [56] data) results in strong performance both indoor 21 and outdoor $2$. However, training on CA-VQA still yields significantly better indoor performance overall 15, suggesting that our high-quality 3D GT is more effective than OSD’s pseudo-annotations. When using CA-VQA + scale aug. 19, we become competitive with OSD $2$ even on outdoor Distance, but still lack on outdoor Width / Height. Combining CA-VQA with OSD 13 results in significant improvements, emphasizing the complementary benefits of the two datasets.
@@ A.1 Spatial Task Categories
    • Viewpoint-Dependent. We consider the spatial relationships left vs. right and in front vs. behind between two objects, as determined from the current camera pose / viewpoint: [^2] Note that we do not consider above vs. below to avoid ambiguity: “above” could either refer to 2D image space (i.e., the 2D bounding box of A is above that of B), or to 3D space, where the latter can be ambiguous as well (i.e., do we just require that the 3D bounding box of A is located higher in terms of vertical dimension, or do we also require that A is located directly above B in terms of horizontal dimensions – the latter might best match with how humans colloquially define “above”).
@@
    • Counting. We always ensure that 0 is an option (i.e., object is not present). We then randomly sample (additional) wrong options among the non-zero integers within $\text{GT}-3,\text{GT}+3]$ (where GT is the correct answer), s.t. the total number of options is 4.
@@ A.2 Depth: Chain-of-Thought (CoT) / Tool-Use
We prepare multi-step CoT responses involving depth for questions within the Binary (only “behind vs. in front”) and Regression (Metric Estimation) categories, as the ground truth answers for those rely on depth information. We also did preliminary experiments with 3D Grounding samples, but found that performance does not improve / even slightly regresses there, so we did not include any such samples in the final dataset. [^3] We hypothesize that 3D grounding is too complex of a task to benefit from the simple depth information provided in the multi-step CoT answers, and that the model might just get confused. We leave a more comprehensive study of how to benefit 3D grounding with CoT for future work.
@@ Appendix D Analysis of Blind Filtering Procedure
Table 11: CA-VQA Blind Filtering Analysis. We study how the improvement from using vision (i.e., comparing a vision-evaluated model vs. a blind-evaluated model) changes after applying the blind filtering strategy outlined in Sec. 3.1, which follows [25]. Our results confirm that after applying our filtering strategy, 1) blind models perform substantially worse, and 2) vision improvements (i.e., the delta between vision and blind models) increase substantially, for both GPT-4/V and $M$-Spatial. This highlights the effectiveness of our blind filtering procedure in ensuring that our CA-VQA benchmark becomes more reliant on vision input (i.e., less susceptible to a strong language prior).
@@
Figure 8: Comparative visualization of axis-aligned vs. oriented 3D bounding boxes, taken from the SpatialRGPT paper [27, Appendix K, Figure 11]. The object dimensions computed from AABBs can differ substantially from those computed from OBBs, depending on the object’s rotation. For sake of illustration, assume that the sofa is 2m wide and 0.8m deep. We then obtain the following altered object dimensions when using an AABB instead of an OBB, at different yaw rotation angles (i.e., considering 7-DOF bounding boxes that are gravity-aligned / parallel to the ground, as in CA-1M / CA-VQA): width $\approx$ 2.1m and depth $\approx$ 1.7m with 30 rotation; width $\approx$ 1.7m and depth $\approx$ 2.1m with 60 rotation; and width = 0.8m and depth = 2m with 90 rotation (i.e., “full” rotation resulting in swapped dimensions).